OpenStudy (anonymous):

Person 1 can do a certain job in twenty-one minutes, and person 2 can do the same job in twenty-eight minutes. When completing the job together, which expression would be used to represent the amount of work done by person 1?

OpenStudy (anonymous):

21 1/21 x/21 21x

OpenStudy (anonymous):

\[ \frac 1{21}+ \frac 1{20}=\frac 1 x \] 1/21 is the portion of the job done by Person 1 in one minute. 1/20 is the portion of the job done by Person 2 in one minute. 1/x is the portion of the job done by in one both in one minute. I think the problem is not worded clearly